an airplane travwl 100km/h then another airplane. In 4 hours the faster airplanes travel the same distance as the slower airplane travel in 6 hours. Find the rate of the faster airplane.
Found 2 solutions by rapaljer, stanbon:
Answer by rapaljer(4671)

Let x = speed of the slower plane
x+100 = speed of the faster plane

Distance = Rate * Time
4(x+100)=6x
4x+400=6x
400=2x
x= 200 mph slower plane
x+100 = 300 mph faster plane

Check: Faster plane (300 mph) in 4 hours travels 1200 miles
Slower plane (200 mph) in 6 hours travels 1200 miles.

R^2

Answer by stanbon(75887)

an airplane travwl 100km/h slower then another airplane. In 4 hours the faster airplanes travel the same distance as the slower airplane travel in 6 hours. Find the rate of the faster airplane.
-------------
Slower plane DATA:
rate = x-100 km/h ; time = 6 hrs; distance = rt = 6(x-100) km
--------------
Faster plane DATA:
rate = x km/h ; time = 4 hrs; distance = rt = 4x km
--------------
Equation:
distance = distance
6(x-100) = 4x
2x = 600
x = 300 km/h (rate of the faster airplane)
x-100 = 200 km/h ( rate of the slower airplane)
=========================
Cheers,
Stan H.